in a bag are 7 pennies, 4 nickels, and 5 dimes. three coins are selected at random. find the probability of each of the following selections.
a.all 3 pennies b.all 3 nickels c.all 3 dimes d.2 pennies,1 dime e.1 penny,1 dime,1 nickel f.1 dime, 2 nickels

Answer by stanbon(48516)

in a bag are 7 pennies, 4 nickels, and 5 dimes. three coins are selected at random. find the probability of each of the following selections.
a.all 3 pennies
7C3/16C3 = 0.0625
---------
b.all 3 nickels
4C3/16C3 = 4/560 = 0.0071
-------------
c.all 3 dimes
5C3/16C3 =
OK. Do you see the picture?
cheers,
Stan H.
d.2 pennies,1 dime e.1 penny,1 dime,1 nickel f.1 dime, 2 nickels